# LLMCheckup: Conversational Examination of Large Language Models via Interpretability Tools

### v1

@@ 5 Evaluation
5.1 Data augmentation
We assess the quality of the generated augmented output based on two key aspects: (1) consistency: the metric represents the proportion of instances where the augmentation process does not lead to a change in the label Li et al. (2023a); Dai et al. (2023); (2) fluency: assesses how well the augmented output aligns with the original data in terms of semantic similarity Ross et al. (2021) measured by SBERT [^7] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Table 1 indicates that Mistral and Llama2 exhibit comparable performance, while Stable Beluga 2 displays substantially higher consistency scores on two tasks, although it may exhibit lower fluency in certain cases. The overall performance on ECQA is relatively lower compared to COVID-Fact. This difference in performance can be attributed to the increased complexity and difficulty of the ECQA task.
@@ 5.2 Parsing
To assess the ability of interpreting user intentions by LLMs, we quantify the performance of each deployed model by calculating the exact match parsing accuracy Talmor et al. (2017); Yu et al. (2018) on a manually created test set, which consists of a total of 119 pairs of user questions and corresponding SQL-like queries. As an additional baseline, we employ the nearest neighbor approach that relies on comparing semantic similarity.
We assess parsing accuracy of our two approaches, GD and MP (§2.1). Table 2 shows that, as model size increases, the parsing accuracy tends to increase and MP demonstrates a notable improvement over GD. Despite Stable Beluga 2 having a larger size compared to 7B models, its parsing performance only marginally surpasses that of Mistral and Llama2. This can be partially attributed to the difficulty of the parsing task [^8] We have a total of 21 operations displayed in Table 6, and many of these offer multiple options. For instance, score operation supports $F_{1}$, precision, recall and accuracy matrices. and the number of demonstrations, as larger models may require a greater number of demonstrations to fully comprehend the context Li et al. (2023b). Parsing results are discussed in more detail in Appendix B.
@@
6 Discussion
We highlight the broad generalizability of LLMCheckup concerning diverse NLP tasks and models. Our approach empowers any auto-regressive LLM to handle dialogue-based explainability including parsing and explanation generation. Aside from the NLP tasks showcased in Section 4, LLMCheckup can be easily expanded to accommodate other tasks not directly related to classification, such as summarization or translation [^9] For other NLP tasks, the predict operation should be replaced by a task-specific operation and all explanation operations can be retained and easily adapted.. In contrast to ConvXAI Shen et al. (2023) and InterroLang Feldhus et al. (2023), which require a fine-tuned model for each specific use case, LLMs used in LLMCheckup possess remarkable zero-/few-shot capabilities Brown et al. (2020), which allows them to effectively handle many tasks without requiring fine-tuning. Although the quality of an explanation could be enhanced with further fine-tuning, LLMCheckup uses model outputs out of the box.
7 Related work
Interfaces for interactive explanations
LIT Tenney et al. (2020) is a GUI-based tool available for analyzing model behaviors across entire datasets. However, LIT has less functionalities in terms of prompting and lower accessibility, e.g. no tutorial and a lower level of integration with HuggingFace. CrossCheck Arendt et al. (2021) exhibits the capability to facilitate quick cross-model comparison and error analysis across various data types, but adapting it for other use cases may need a substantial amount of code modification and customization. XMD’s Lee et al. (2023) primary purpose is model debugging, but it shares similarities in the focus on feature attributions, visualization of single instances and user feedback options. It is, however, limited to feature attribution explanations and smaller, efficiently retrainable models. IFAN Mosca et al. (2023) enables real-time explanation-based interaction with NLP models, but is limited to the sequence-to-class format, restricting its applicability to other tasks. Furthermore, IFAN offers a limited set of explainability methods, consisting of only one local and one global method.
Dialogue-based systems for interpretability
Carneiro et al. (2021) point out that conversational interfaces have the potential to greatly enhance the transparency and the level of trust that human decision-makers place in them. According to Zhang et al.’s user studies, delivering explanations in a conversational manner can improve users’ understanding, satisfaction, and acceptance. Jacovi et al. (2023) emphasizes the necessity of interactive interrogation in order to build understandable explanation narratives. ConvXAI Shen et al. (2023), TalkToModel Slack et al. (2023), InterroLang Feldhus et al. (2023) and Brachman et al. (2023) share some similarities with our system, but are more complex in their setup and consider fewer explainability methods. Additionally, they might overrely on external LMs to explain the deployed model’s behavior, whereas LLMCheckup places a strong emphasis on self-explanation, which is crucial for faithfulness. Finally, LLMCheckup uses auto-regressive models, as they have become increasingly dominant in various NLP applications nowadays. In iSee Wijekoon et al. (2023), a chatbot adapts explanations to the user’s persona. Explanation strategies and user experiences can be captured and shared. The system is currently still in development, but shares many functionalities, except that LLMs are not in their scope yet.
8 Conclusion
We present the interpretability tool LLMCheckup, designed as a dialogue-based system. LLMCheckup can provide explanations in a conversation with the user facilitated by any auto-regressive LLM. By consolidating parsing, downstream task prediction, explanation generation and response generation within a unified framework, LLMCheckup streamlines the interpretability process and serves as a baseline for future investigation.
Future work includes exploring RAG models Lewis et al. (2020) combined with explainability, as currently LLMCheckup relies on search engines for external information retrieval. We also want to add multi-modal models, so converting image or audio input to texts would no longer be necessary, but the current state of interpretability on such models lags behind unimodal approaches Liang et al. (2023). Integrating our system into HuggingChat [^10] https://huggingface.co/chat/ would further increase the visibility and accessibility through the web.
Limitations
English is the main language of the current system. Multilinguality is not supported, as both the interface, the responses, tutorial and the explained models are monolingual. While it would be possible to adapt it to other languages by translating interface texts and prompts and using a model trained on data in another target language or multiple ones, it remains to be seen to which extent multilingual LLMs can do quadruple duty as well as the current model does for English.
In LLMCheckup, users have the flexibility to input data in different modalities, including images and audio. However, for audio and images, LLMCheckup will convert the audio content and texts contained within the images into textual format for further processing and analysis. Besides, the explanations and responses generated by our system are currently limited to the text format – apart from the heatmap visualization of feature attribution explanations.
Appendix A Explanation examples
User LLMCheckup
Hi! What do you predict for instance with id 2435?
10 most important features for it regarding attention?
And starting it how would a new instance look like?
Can you explain it in natural language?
User LLMCheckup
How would you flip the prediction for it?
Is it possible to retrieve an example that is similar to it?
Appendix B Multi-prompt parsing
As a first step, MP queries the model about the main operation (e.g., feature attribution, counterfactual generation, dataset statistics etc., see the full list of the available operations in Appendix D). Next, depending on the chosen operation, MP selects the operation-specific prompts with 2-7 demonstrations [^11] The number of demonstrations depends on the difficulty of operation, e.g., how many attributes it may have. (user query and parsed outputs examples) to generate the full parses that may include several attributes. E.g., for the user input "What are the feature attributions for id 42 based on the integrated gradients?" we start by generating nlpattribute and then augment the parse with the second prompt and transform it into filter id 42 and nlpattribute integrated_gradient.
Since the output of the model is not constrained compared to GD, in the MP setting, we need to check whether the model’s output matches any of the available operations and if there is no exact match we employ SBERT and compare the embedding similarity between the user input and the operations of LLMCheckup (e.g., sometimes the model outputs synonyms but we need to match the operation names exactly to ensure the correct parsing). Additionally, we implement some checks to avoid possible hallucinations, e.g., if the model outputs an id that is not present in the user input we remove it from the parsed text.
@@ Appendix B Multi-prompt parsing
Table 5 summarizes our parsing evaluation results for different models with different number of ’max_new_tokens‘ for generation specified in the configuration file. We found that this parameter is very important for the performance of the MP parser. Pythia-2.8B, Mistral-7B and Stable Beluga 2 achieve better scores with less tokens to generate (10) and Llama2-7B shows better performance with larger number of tokens (20). The only model that is unaffected by the ’max_new_tokens‘ parameter is Falcon-1B which also has the smallest size. The majority of the models perform better when allowed to generate only brief texts and if they are given more tokens to generate they tend to hallucinate and detract from the actual parsing task.
@@
Originally, we treated all models in the same way during parsing, however, after looking at some generated outputs we realized that Falcon-1B and Pythia-2.8B are not good at extracting ids and often can only recognize the main operation. Hence, for these models we have an additional step that extracts a potential id from the user input and adds it to the parsed operation.
As expected, larger models tend to perform better than the ones with less parameters. However, we also found that the quantized Llama model outperforms its full (non-quantized) version on the parsing task.
Appendix C Follow-up suggestions
We group the suggestions into several categories that include metadata operations (e.g., dataset statistics, model types etc.), prediction-related operations (e.g., predict, count or show mistakes), explanation-based operations (e.g., attributions for top $k$, attention scores and integrated gradients or free-text rationale), NLU (similarity and keywords) and input perturbations (counterfactuals and data augmentation).
The user always has an option to decline the suggestion and ask something different. We check whether the user agrees to the LLMCheckup suggestions by computing the similarity scores between the input and the confirm/disconfirm templates with SBERT.
Appendix D Supported operations
We define an operation as a mapping between (a part of) an SQL-type query and a function executed by LLMCheckup. Table 6 enumerates all operations facilitated by LLMCheckup, including all explainability methods mentioned in §3 and other supplementary operations.
Filt. filter(id) Access single instance by its ID
includes(token) Filter instances by token occurrence
Prediction $*{}^{*}$ Get the prediction of the given instance
randompredict(number) Precompute a subset of instances at random
mistakes(subset) Count number of wrongly predicted instances
Data show(list) Showcase a list of instance
countdata(list) Count number of instances within the given list
label(dataset) Describe the label distribution across the dataset
Meta data() Information related to the dataset
model() Metadata of the model
About function() Inform the functionality of the system
self() Self-introduction
qatutorial(op_name) Provide explanation for the given operation
Logic and(op1, op2) Concatenation of multiple operations
or(op1, op2) Selection of multiple filters
Explanation nlpattribute(inst., topk, $*{}^{*}$ Provide feature attribution scores
$*{}^{*}$ Explain the output in natural language
$*{}^{*}$ Generate counterfactuals
similarity(instance, $*{}^{*}$ Output top $k$ similar instances in the dataset
$*{}^{*}$ Augment the input text
Table 6: Operations marked with ($*{}^{*}$) support the use of custom inputs.
Appendix E Details on feature attribution
@@
Appendix F Interface
Figure 3 display the user interface of LLMCheckup. Within the chat window, a user’s question requesting the explanation for instance with ID 4180 in natural language is shown, followed by a response that includes the corresponding rationale. On the right side of the window, there is a prompt modification section where the Optimization by Prompting (OPRO) strategy is chosen for rationalization. The icons associated with each strategy describe them in detail when they are hovered over, including the corresponding prompts that can be appended after the system prompt we have designed.
Figure 3: LLMCheckup interface with initial welcome message, generated free-text rationale and sample generator buttons. Expert level of knowledge in XAI is selected and Optimization by Prompting (OPRO) strategy is used. For an example dialogue, see Table 3 and Table 4.
Appendix G QA tutorial
Figure 4 shows tutorials for data augmentation with different level of expertise in XAI.
@@ Appendix G QA tutorial
Appendix H External information retrieval

### v2

@@ 5 Evaluation
We conducted evaluations for parsing and data augmentation with LLMs using automated evaluation metrics [^9] Note that our evaluation does not involve any user study, as that aspect is considered as future work and falls outside the scope of our initial focus on engineering.. Among all the supported methods presented in Table 1, we chose data augmentation as a representative operation to evaluate the performance of different LLMs.
@@ 5.1 Parsing
To assess the ability of interpreting user intents by LLMs, we quantify the performance of each deployed model by calculating the exact match parsing accuracy Talmor et al. (2017); Yu et al. (2018) on a manually created test set, which consists of a total of 119 pairs of user questions and corresponding SQL-like queries. As an additional baseline, we employ the nearest neighbor approach that relies on comparing semantic similarity.
We assess parsing accuracy of our two approaches, GD and MP (§2.2). Table 2 shows that, as model size increases, the parsing accuracy tends to increase and MP demonstrates a notable improvement over GD. Despite Stable Beluga 2 having a larger size compared to 7B models, its parsing performance only marginally surpasses that of Mistral and Llama2. This can be partially attributed to the difficulty of the parsing task [^10] We have a total of 21 LLMCheckup operations displayed in Table 1 (excluding the logic operations), and many of these offer multiple options. For instance, score operation supports $F_{1}$, precision, recall and accuracy matrices. and the number of demonstrations, as larger models may require a greater number of demonstrations to fully comprehend the context Li et al. (2023b).
@@
Table 3 summarizes our parsing evaluation results for different models with different number of ’max_new_tokens‘ for generation. Llama-based models showed better performance with more tokens to generate compared to the rest of the models. After looking at some generated outputs we realized that Falcon-1B and Pythia-2.8B are not good at extracting ids and often can only recognize the main LLMCheckup operation. Hence, for these two models we have an additional step that extracts a potential ID from the user input and adds it to the parsed operation. As expected, larger models tend to perform better than the ones with fewer parameters. However, we also found that the quantized Llama model outperforms its full (non-quantized) version on the parsing task.
5.2 Data augmentation
Dataset COVID-Fact ECQA
Mistral 7B 0.66 0.88 0.96 0.50 0.76
Llama2 7B 0.65 0.88 0.94 0.50 0.76
Stable Beluga 2 70B 1.00 0.85 0.96 1.00 0.73
We assess the quality of the generated augmented output based on two key aspects: (1) consistency: the metric represents the proportion of instances where the augmentation process does not lead to a change in the label before and after the augmentation Li et al. (2023a); Dai et al. (2023); (2) fluency: assesses how well the augmented output aligns with the original data in terms of semantic similarity Ross et al. (2021) measured by SBERT. Table 4 indicates that Mistral and Llama2 exhibit comparable performance, while Stable Beluga 2 displays substantially higher consistency scores on two tasks, although it may exhibit lower fluency in certain cases. The overall performance on ECQA is relatively low compared to COVID-Fact. This difference in performance can be attributed to the increased complexity of the ECQA task. Our primary focus is to compare the performance of different LMs (Table 4), rather than aiming for state-of-the-art results on both downstream tasks or demonstrating perfect fluency and consistency [^11] Creating gold data is out of scope for this work, because it involves costly human annotations. For the lack of gold data, we have intentionally omitted providing a baseline..
6 Discussion
In contrast to previous dialogue-based XAI frameworks ConvXAI Shen et al. (2023) and InterroLang Feldhus et al. (2023), which require a fine-tuned model for each specific use case, LLMs used in LLMCheckup possess remarkable zero-/few-shot capabilities Brown et al. (2020) for effectively handling many tasks without requiring fine-tuning. Although the quality of an explanation could be enhanced with further fine-tuning, LLMCheckup uses model outputs out of the box.
Our empirical results underline the feasibility of conversational interpretability and the usefulness of LLMCheckup for future studies, especially human evaluation. We focus on the ground work in terms of engineering, implementation and user interface, for connecting the human with the model. This provides user studies Wang et al. (2019); Feldhus et al. (2023); Zhang et al. (2023) in the future with a head start, s.t. they can spend more time on conducting their study. We see evaluation measures for differences between users’ mental models and model behavior and objective metrics beyond simulatability as the most important gaps to fill.
7 Related work
Interfaces for interactive explanations
LIT Tenney et al. (2020) is a GUI-based tool available for analyzing model behaviors across entire datasets. However, LIT has less functionalities in terms of prompting and lower accessibility, e.g. no tutorial and a lower level of integration with HuggingFace. CrossCheck Arendt et al. (2021) exhibits the capability to facilitate quick cross-model comparison and error analysis across various data types, but adapting it for other use cases needs substantial code modification and customization. XMD’s Lee et al. (2023) primary purpose is model debugging, but it shares similarities in the focus on feature attributions, visualization of single instances and user feedback options. It is, however, limited to feature attribution explanations and smaller, efficiently retrainable models. IFAN Mosca et al. (2023) enables real-time explanation-based interaction with NLP models, but is limited to the sequence-to-class format, restricting its applicability to other tasks and it offers only a limited set of explainability methods.
Dialogue-based systems for interpretability
Carneiro et al. (2021) point out that conversational interfaces have the potential to greatly enhance the transparency and the level of trust that human decision-makers place in them. According to Zhang et al.’s (2023) user studies, delivering explanations in a conversational manner can improve users’ understanding, satisfaction, and acceptance. Jacovi et al. (2023) emphasizes the necessity of interactive interrogation in order to build understandable explanation narratives. ConvXAI Shen et al. (2023), TalkToModel Slack et al. (2023), InterroLang Feldhus et al. (2023) and Brachman et al. (2023) share some similarities with our framework, but are more complex in their setup and consider fewer explainability methods. Additionally, they might overrely on external LMs to explain the deployed LM’s behavior, whereas LLMCheckup places a strong emphasis on self-explanation, which is crucial for faithfulness. Finally, LLMCheckup uses auto-regressive models, as they have become increasingly dominant in various NLP applications nowadays. In iSee Wijekoon et al. (2023), a chatbot adapts explanations to the user’s persona, but they do not consider LLMs.
8 Conclusion
We present the interpretability tool LLMCheckup, designed as a dialogue-based system. LLMCheckup can provide explanations in a conversation with the user facilitated by any auto-regressive LLM. By consolidating parsing, downstream task prediction, explanation generation and response generation within a unified framework, LLMCheckup streamlines the interpretability process without switching between different LMs, modules or libraries and serves as a baseline for future investigation.
Future work includes exploring RAG models Lewis et al. (2020) combined with explainability, as currently LLMCheckup relies on search engines for external information retrieval. We also want to add multi-modal models, so that converting image or audio input to texts would no longer be necessary, but the current state of interpretability on such models lags behind unimodal approaches Liang et al. (2023). Integrating our framework into HuggingChat [^12] https://huggingface.co/chat/ would further increase the visibility and accessibility through the web.
Limitations
English is the main language of the current framework. Multilinguality is not supported, as both the interface, the responses, tutorial and the explained models are monolingual. While it would be possible to adapt it to other languages by translating interface texts and prompts and using a model trained on data in another target language or multiple ones, it remains to be seen to which extent multilingual LLMs can do quadruple duty as well as the current model does for English.
In LLMCheckup, users have the flexibility to input data in different modalities, including images and audio. However, for audio and images, LLMCheckup will convert the audio content and texts contained within the images into textual format for further processing and analysis. Besides, the explanations and responses generated by our framework are currently limited to the text format – apart from the heatmap visualization of feature attribution explanations.
Acknowledgement
We thank the anonymous reviewers of the NAACL HCI+NLP Workshop for their constructive feedback on our paper. This work has been supported by the German Federal Ministry of Education and Research as part of the project XAINES (01IW20005).
Appendix A Supported operations in LLMCheckup
Table 1 lists all operations supported by LLMCheckup. Operations other than those related to explanation (Table 5, Table 6) are considered supplementary and are responsible for providing statistics and meta-information about data, model or LLMCheckup to make it more user-friendly. For instance, predict operation enables users to receive predictions and serves as an initial step for starting an explanatory dialogue; data operation can offer meta-information about the dataset, thereby sharing essential background knowledge with the users, when they start a new dialogue.
Appendix B Explanation examples
User LLMCheckup
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
User LLMCheckup
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image]
Appendix C Details on feature attribution
@@
Appendix D QA tutorial
Figure 4 shows tutorials for data augmentation with different levels of expertise in XAI.
@@ Appendix D QA tutorial
Appendix E External information retrieval
@@
Appendix F Suggestion of follow-up questions
The suggestion mode can provide follow-up questions for metadata operations (e.g., dataset statistics, model types etc.), prediction-related operations (e.g., predict, count or show mistakes), explanation-based operations (e.g., attributions for top $k$, attention scores and integrated gradients or free-text rationale), NLU (similarity and keywords) and input perturbations (counterfactuals and data augmentation). These categories are also summarized in Table 1.
The user always has an option to decline a suggestion and ask something different. We check whether the user agrees with the LLMCheckup suggestions by computing the similarity scores between the input and the confirm/disconfirm templates with SBERT.
Additionally, for each generated suggestion we check whether it already appears in the dialogue history to make sure that the user does not receive repetitive suggestions for the operations that have already been performed. E.g., if the user inquires about the counterfactual operation and the model explains how it works, LLMCheckup will store this information and will not suggest explaining counterfactuals again.